OpenStudy (anonymous):

Which of the following equations is the result of completing the square on x^2 - 6x - 9 = 0?

OpenStudy (anonymous):

(x - 3)2= 0******8 (x - 3)2= 9 (x - 3)2= 18 the one with the ***** is the one I think is right im I right

Parth (parthkohli):

\[(x^2 - 6x + 9) - 18=0 \quad \Rightarrow \quad (x - 3)^2 = 18\]

OpenStudy (goformit100):

I will help if you Unblock me.

Parth (parthkohli):

\[x - 3 = \pm \sqrt{18}\]Can you finish?

OpenStudy (anonymous):

x^2-6x+9=18